OpenStudy (anonymous):

The radius of the base of a cylinder is 28 cm and its height is 48 cm. Find the surface area of the cylinder is terms of pi.

OpenStudy (anonymous):

2π(28)(28+48) = 56π(76) = 4256π

OpenStudy (anonymous):

area of the bases \[2\pi r ^{2}=2\pi * 784=1568\pi\] circumference of base \[2\pi r = 56\pi\] extension of the height \[2\pi r*h=56\pi*48=2688\pi\] bases plus none base surface is total surface \[1568\pi + 2688\pi = 4256\pi\]

OpenStudy (anonymous):

Surface Area of cylinder = 2πr (r+h) where r = radius of base h = height of cylinder